the perimeter of a rectangle is 114 feet. its length is three more than twice its width. can you help me solve this please ?

Answer by edjones(8007)

2L+2w=114
L=2w+3
.
2(2w+3)+2w=114
4w+6+2w=114
6w=108
w=18'
L=39'
.
Ed